OpenStudy (anonymous):

Hey guys. Will you help me find the solution to this function? g(x)=4x2+8x+2 Fan/medal for help and answer. :)

OpenStudy (anonymous):

I can help but no medal please

OpenStudy (anonymous):

Okay. :)

OpenStudy (anonymous):

but then what will you give?

OpenStudy (anonymous):

I'll fan you?

OpenStudy (anonymous):

Or I can try to help on anything you may need help with. I'm good at English and Chemistry.

OpenStudy (anonymous):

K sounds good g(x)=0

OpenStudy (anonymous):

How'd you get that?

OpenStudy (anonymous):

4x2+8x+2 = 0 so 4x^2+4x+2x+2=0 x = -1 or x =-1/2

OpenStudy (anonymous):

wait I made a mistake

OpenStudy (***[isuru]***):

hate to say this @thisSucks ,,, but u r wrong!!!

OpenStudy (***[isuru]***):

u can't factorize that one ....

OpenStudy (anonymous):

you could also use the quadratic formula to solve this

OpenStudy (***[isuru]***):

u have to use either completing square or the formula

OpenStudy (***[isuru]***):

but before that .... is g(x) = 0 ?

OpenStudy (anonymous):

Yeah SURU is right

OpenStudy (anonymous):

you have to use the formula

OpenStudy (***[isuru]***):

u can't solve g(x) , if g(x) is just an expression...

OpenStudy (anonymous):

Completing the square is actually the name of the lesson. All I know is that it will have two irriational solutions due to the rules of discrimnants.

OpenStudy (***[isuru]***):

so.... go on with completing square....

OpenStudy (anonymous):

Yeah so you can apply the formula there. Sorry I made the mistake :(

OpenStudy (***[isuru]***):

let's assume that g(x) = 0 ... Otherwise u won't get any answers...

OpenStudy (anonymous):

(2x+2)^2 -2 =0

OpenStudy (***[isuru]***):

first step.... \[g(x) =4x ^{2} +8x +2\]

OpenStudy (***[isuru]***):

take 4 as common factor... out of the expression

OpenStudy (anonymous):

Okay. But what happens with the two?

OpenStudy (***[isuru]***):

\[g(x) = 4( x ^{2} + 2x + \frac{ 1 }{ 2 })\]

OpenStudy (***[isuru]***):

now... wt is the next step... in ur lesson ?

OpenStudy (***[isuru]***):

divide the coefficient of x by 2 and write x and that value of the coefficient as a complete square...

OpenStudy (***[isuru]***):